# The Manga Whisperer: Automatically Generating Transcriptions for Comics

### v1

@@ 3.2 Training
Implementation: The input image is resized such that the shorter side is 800800px and the longer side is resized appropriately to preserve the aspect ratio but capped at 13331333px. It is then projected to a spatial feature map by the backbone $\Phi_{b}(\cdot)$, which has the ResNet-50 [10] architecture. The dimensions of the spatial feature map depend on the resolution of the input image (e.g. for a resized $1100\times 800$px image, the spatial map is $35\times 25$). The spatial dimension of this feature map is flattened and the channel dimension is projected to $256$ before passing it to the transformer. The transformer encoder-decoder $\Phi_{e}(\cdot),\Phi_{d}(\cdot)$ have 6 layers each, with the hidden dimension of $256$, and $8$ attention heads. The number of [OBJ] decoder queries is $N=300$. $\Phi_{b}(\cdot),\Phi_{e}(\cdot),\Phi_{d}(\cdot),\Phi_{box}(\cdot),\Phi_{cls}(\cdot)$, are initialised using Conditional-DETR [29] weights. The crop-embedding module $\Phi_{crop}(\cdot)$ has 12 layers, with the hidden dimension of 768 and 12 attention heads, and is initialised using MAE [11] encoder weights. The input to $\Phi_{crop}(\cdot)$ is resized to $224\times 224$. The character matching and speaker association modules $\Phi_{c2c},\Phi_{t2c}$ are both 3-layered MLPs initialised randomly. We set the character matching threshold $\tau=0.7$. Our training objective for detection is the same as in [29]. We further apply Binary Cross Entropy loss to the outputs of our character matching and speaker association heads. Additionally, we apply Supervised Contrastive Loss [17] to the per-page embeddings from the crop-embedding module. We trained our model end-to-end, on $2\timesA$40 GPUs using AdamW [25] optimizer with both learning rate and weight decay of 0.0001, and effective batch size of 16.
@@ 4 Transcript Generation
Panel Ordering: Previous methods [18, 13] use “cuts” to recursively split the panels into horizontal and vertical partitions. The idea is to construct a tree by recursively splitting the panels using (1) a horizontal line, (2) a vertical line, in that order of priority, and then traverse the tree in a specific order (top panels before bottom for horizontal cuts, right panels before left for vertical cuts) to get the reading order. While this method works in most cases, it fails in the case of overlapping panels, where there is no clean way to “cut” the page. We propose an improved algorithm to sort the panels in their reading order that overcomes this issue. We very briefly describe our approach below and provide more details in the supplementary.
@@
OCR: While performing OCR is not the primary goal of this work, we train an OCR model for completeness. Specifically, we finetune the TrOCR [20] model on custom synthetic data generated using the pipeline described in [4]. We provide more details in the supplementary.
@@
Figure 5: Text to speaker predictions generated by the Magi model. Each predicted text box is connected to a predicted character box using a line. The opacity of the line reflects the confidence of the model (the darker the line, the more confident the model is). Each predicted character box has a number at its centre based on the clustering predictions. We also show the final generated transcript. Note that all the dialogues are in the correct reading order. For text to speaker predictions that have a low confidence score (<0.4<0.4) we replace the predicted speaker with $\langle?\rangle$ in the generated transcript and let the reader infer it from context. Best viewed digitally.
@@ 5.1 Dataset overview
PopManga: A new dataset of 57000+ manga pages from $80+$ different series that are considered to be in the list the top manga of all time [32]. These manga are typically known for their complex storytelling and intricate art style, thus making them very challenging. We split PopManga into 3 subsets. The Dev subset is used for training and validation, and the remaining two subsets—Test-S (test seen), Test-U (test unseen)—are used for testing such that Test-S consists of 30 chapters from 15 series that the model sees during training while Test-U consists of 20 chapters from 10 series that the model never sees during training.
We intend the Test-S and Test-U images to serve as a public benchmark for evaluation. For this reason, they are from chapters that are available freely, publicly and officially on Manga Plus by Shueisha [28], thus enabling future academic research and comparisons. A comprehensive list of manga name and chapters in Dev, Test-S, Test-U is available in the supplementary.
@@ 5.2 Dataset curation and annotations
Annotation: We get human annotators to label text boxes, character boxes, character-character association and text-character association using a modified version of the LISA annotation tool [9]. For panels, we first finetune DETR [5] to predict panels using Manga109 and then use it to automatically generate pseudo panel annotations for for the Dev subset of PopManga (we found these generated panels to be of satisfactory quality, thus reducing annotation cost). In the supplementary material we include more information on the annotation process.
@@ 6.2 Evaluation Metrics
Magi (crop only) [Ours] 0.4342 0.6734 0.8808 0.7452 0.8099 0.7655
Magi [Ours] 0.6282 0.8316 0.9210 0.8238 0.8737 0.8373
@@
Magi (crop only) [Ours] 0.4387 0.7022 0.8801 0.7445 0.8059 0.7626
Magi [Ours] 0.6282 0.8389 0.9191 0.8218 0.8692 0.8336
@@
Magi (crop only) [Ours] 0.3182 0.6078 0.8233 0.6370 0.7099 0.6628
Magi [Ours] 0.5088 0.7709 0.9084 0.8013 0.8469 0.8152
@@
Magi (crop only) [Ours] 0.5132 0.7322 0.9052 0.7887 0.8418 0.8050
Magi [Ours] 0.6116 0.8068 0.9295 0.8385 0.8824 0.8503
@@ C.1 Data splits
Note that the Test-S chapters are from the “seen” manga i.e. the model has seen other chapters from this manga during training, whereas the Test-U chapters are from the “unseen” manga i.e. the model has never seen any chapter from this manga during training. In total there over 2000 chapters in Dev, 30 chapters in Test-S and 20 chapters in Test-U.
@@ D.2 Implementation of metrics
    While computing AMI and NMI, we do not assume prior knowledge regarding the number of unique identities present on the page. This is because we wish to evaluate the model’s ability to form decisive clusters without knowing the number of clusters apriori. To facilitate this, we compute a character-character similarity matrix for each page and threshold this matrix to give us binary values for whether two characters are the same or not (thus, giving us clusters). The AMI and NMI results reported in Table 3 for baselines are by picking the best threshold for each method for each test set (where “best” corresponds to the threshold value with maximum AMI results). The threshold for Magi is always $0.7$.

### v3

@@ 3.2 Training
Implementation: The input image is resized such that the shorter side is $800$px and the longer side is resized appropriately to preserve the aspect ratio but capped at $1333$px. It is then projected to a spatial feature map by the backbone $\Phi_{b}(\cdot)$, which has the ResNet-50 [10] architecture. The dimensions of the spatial feature map depend on the resolution of the input image (e.g. for a resized $1100\times 800$px image, the spatial map is $35\times 25$). The spatial dimension of this feature map is flattened and the channel dimension is projected to $256$ before passing it to the transformer. The transformer encoder-decoder $\Phi_{e}(\cdot),\Phi_{d}(\cdot)$ have 6 layers each, with the hidden dimension of $256$, and $8$ attention heads. The number of [OBJ] decoder queries is $N=300$. $\Phi_{b}(\cdot),\Phi_{e}(\cdot),\Phi_{d}(\cdot),\Phi_{box}(\cdot),\Phi_{cls}(\cdot)$, are initialised using Conditional-DETR [29] weights. The crop-embedding module $\Phi_{crop}(\cdot)$ has 12 layers, with the hidden dimension of 768 and 12 attention heads, and is initialised using MAE [11] encoder weights. The input to $\Phi_{crop}(\cdot)$ is resized to $224\times 224$. The character matching and speaker association modules $\Phi_{c2c},\Phi_{t2c}$ are both 3-layered MLPs initialised randomly. We set the character matching threshold $\tau=0.65$. Our training objective for detection is the same as in [29]. We further apply Binary Cross Entropy loss to the outputs of our character matching and speaker association heads. Additionally, we apply Supervised Contrastive Loss [17] to the per-page embeddings from the crop-embedding module. We trained our model end-to-end, on 2×2$\times$2 ×A40 GPUs using AdamW [25] optimizer with both learning rate and weight decay of 0.0001, and effective batch size of 16.
@@ 4 Transcript Generation
Panel Ordering: Previous methods [18, 13] use “cuts” to recursively split the panels into horizontal and vertical partitions. The idea is to construct a tree by recursively splitting the panels using (1) a horizontal line, (2) a vertical line, in that order of priority, and then traverse the tree in a specific order (top panels before bottom for horizontal cuts, right panels before left for vertical cuts) to get the reading order. While this method works in most cases, it fails in the case of overlapping panels, where there is no clean way to “cut” the page. We propose an improved algorithm to sort the panels in their reading order that overcomes this issue. We very briefly describe our approach below and provide more details in the appendix.
@@
OCR: While performing OCR is not the primary goal of this work, we train an OCR model for completeness. Specifically, we finetune the TrOCR [20] model on custom synthetic data generated using the pipeline described in [4]. We provide more details in the appendix.
@@
Figure 5: Text to speaker predictions generated by the Magi model. Each predicted text box is connected to a predicted character box using a line. The opacity of the line reflects the confidence of the model (the darker the line, the more confident the model is). Each predicted character box has a number at its centre based on the clustering predictions. We also show the final generated transcript. Note that all the dialogues are in the correct reading order. For text to speaker predictions that have a low confidence score ($<0.4$) we replace the predicted speaker with ?delimited-⟨⟩$\langle$?$\rangle$ ? ⟩ in the generated transcript and let the reader infer it from context. Best viewed digitally.
@@ 5.1 Dataset overview
PopManga: A new dataset of 57000+ manga pages from $80+$ different series that are considered to be in the list the top manga of all time [32]. These manga are typically known for their complex storytelling and intricate art style, thus making them very challenging. We split PopManga into 3 subsets. The Dev subset is used for training and validation, and the remaining two subsets—Test-S (test seen), Test-U (test unseen)—are used for testing such that Test-S consists of 30 chapters from 15 series, where other chapters from these series are seen during training, while Test-U consists of 20 chapters from 10 series that the model never sees during training.
We intend the Test-S and Test-U images to serve as a public benchmark for evaluation. For this reason, they are from chapters that are available freely, publicly and officially on Manga Plus by Shueisha [28], thus enabling future academic research and comparisons. A comprehensive list of manga name and chapters in Dev, Test-S, Test-U is available in the appendix.
@@ 5.2 Dataset curation and annotations
Annotation: We get human annotators to label text boxes, character boxes, character-character association and text-character association using a modified version of the LISA annotation tool [9]. For panels, we first finetune DETR [5] to predict panels using Manga109 and then use it to automatically generate pseudo panel annotations for the Dev subset of PopManga (we found these generated panels to be of satisfactory quality, thus reducing annotation cost). In the appendix we include more information on the annotation process.
@@ 6.2 Evaluation Metrics
Magi (crop only) [Ours] 0.4892 0.7178 0.9008 0.7840 0.8423 0.8008
Magi [Ours] 0.6574 0.8501 0.9312 0.8439 0.8884 0.8555
@@
Magi (crop only) [Ours] 0.4862 0.7326 0.9061 0.7926 0.8477 0.8076
Magi [Ours] 0.6527 0.8503 0.9347 0.8557 0.8936 0.8656
@@
Magi (crop only) [Ours] 0.3429 0.6078 0.8381 0.6653 0.7341 0.6896
Magi [Ours] 0.6254 0.7843 0.9148 0.8151 0.8558 0.8279
@@
Magi (crop only) [Ours] 0.5690 0.7694 0.9237 0.8259 0.8721 0.8389
Magi [Ours] 0.6345 0.8202 0.9383 0.8567 0.8966 0.8667
@@ C.1 Data splits
Note that the Test-S chapters are from the “seen” manga i.e. the model has seen other chapters from this manga during training, whereas the Test-U chapters are from the “unseen” manga i.e. the model has never seen any chapter from this manga during training. In total there are over 2000 chapters in Dev, 30 chapters in Test-S and 20 chapters in Test-U.
@@ D.2 Implementation of metrics
    While computing AMI and NMI, we do not assume prior knowledge regarding the number of unique identities present on the page. This is because we wish to evaluate the model’s ability to form decisive clusters without knowing the number of clusters apriori. To facilitate this, we compute a character-character similarity matrix for each page and threshold this matrix to give us binary values for whether two characters are the same or not (thus, giving us clusters). The AMI and NMI results reported in Table 3 for baselines are by picking the best threshold for each method for each test set (where “best” corresponds to the threshold value with maximum AMI results). The threshold for Magi is always $0.65$.